# PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection

### v5

@@ 3.4 Dual Project Constraint
3.5 Objective Function
The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$. To quantify the dissimilarity between these representations, we utilize a comparative loss function based on Kullback–Leibler divergence (KL divergence), termed as Inter-Intra Discrepancy. Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both. The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
@@ 3.6 Anomaly Score
Table 1: Comprehensive comparative results on four datasets. The P, R and F1 are the precision, recall and F1-score. All results are in %, the best in Bold, and the second in underlined.
Dataset MSL SMAP PSM SWaT Average F1
Metric P R F1 P R F1 P R F1 P R F1
OCSVM [19] 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
IForest [37] 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
LOF [8] 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
U-Time [38] 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
ITAD [39] 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
DAGMM [9] 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
VAR [12] 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
MMPCACD [40] 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
CL-MPPCA [41] 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
TS-CP2 [42] 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
BeatGAN [43] 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
LSTM-VAE [6] 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
BOCPD [44] 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
Deep-SVDD [45] 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
LSTM [46] 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
OmniAnomaly [7] 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
InterFusion [47] 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
THOC [48] 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
AnomalyTrans [27] 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
DCdetector [10] 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
TimesNet [14] 89.54 75.36 81.84 90.14 56.40 69.39 98.51 96.20 97.34 90.75 95.40 93.02 85.40
PatchTST [49] 88.34 70.96 78.70 90.64 55.46 68.82 98.84 93.47 96.08 91.10 80.94 85.72 82.33
GPT2-Adapter [15] 82.00 82.91 82.45 90.60 60.95 72.88 98.62 95.68 97.13 92.20 96.34 94.23 86.67
@@
PatchAD (Ours) 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
4 Experiments
4.1 Setups and Evaluation Metrics
We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP (3) PSM (4) SMD (5) SWaT (6) WADI (7) NIPS-TS-SWAN (8) NIPSTS-GECCO (9) UCR. Details about these datasets are outlined in Table 4 in Appendix. Also, implementation and pseudo-code Algorithm 1 details in Appendix.
We compared our model with several benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; (8) Large language model; (9) Diffusion model; See details in Table 1 and Table 5, or in Appendix B.2.
Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as AUC, accuracy, precision, recall, F1-score with point adjustment (F1), and original F1-score (F1_ori) [27]. Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall [51], along with Volume under the surface (VUS) [52]. Although the F1-score is a widely used metric, it does not effectively evaluate anomalous events. Therefore, we introduce the following two new metrics. Affiliation precision and recall are evaluation metrics based on the proximity of predicted events to the true labels. VUS considers including anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve. Employing various methodologies allows for a more comprehensive evaluation of our model. To compare with other methods fairly, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
4.2 Comparison Results
Data Source Clarification: Due to substantial device variations, to ensure a fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes. The results of AnomalyTrans and DCdetector in Table 1 and Table 5 are our reproductions.
From Table 1, PatchAD achieves the best average-level F1 performance. As far as we know, there is ongoing controversy regarding the evaluation of time series anomaly detection. However, precision, recall, and F1-score remain the most widely used evaluation metrics. To comprehensively assess the performance of our approach, additional metrics such as affiliation precision/recall and VUS have been introduced.
In Table 5 in Appendix, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types. Compared to other baseline methods, PatchAD consistently achieves SOTA results on these two datasets. Particularly noteworthy is the 9.3% increase in F1-score over the DCdetector and a 24.9% increase over the Anomaly Transformer on the NIPS-TS-GECCO dataset. Given the remarkable performance of AnomalyTrans and DCdetector surpassing other baselines significantly, we conducted a comprehensive comparison of PatchAD with them across all datasets using multiple metrics, as depicted in Table 7 in Appendix. Our proposed method outperforms or at least remains competitive compared to these two SOTA methods. PatchAD attains superior results across multiple datasets and multiple metrics. Despite PatchAD’s relatively simple network structure, its performance remains competitive. Table 2 demonstrates that our approach outperforms previous methods on univariate dataset UCR. This shows that our method can also fetch state-of-the-art performance on univariate datasets.
Table 2: Results on UCR dataset.
Metric Acc P R F1
AnomalyTrans 99.49 60.41 100 73.08
DCdetector 99.51 61.62 100 74.05
PatchAD (Ours) 99.52 72.72 100 84.00
Figure 2 depicts the average performance of PatchAD compared to various baselines across multiple datasets (WADI, SWAT, PSM, NIPS-TS-SWAN). PatchAD has demonstrated a remarkable improvement in the F1 score, with a notable increase of 50.5%. Additionally, there has been a 7.8% improvement in the Aff-F1 score and a commendable 10.0% boost in the AUC.
@@ 4.2 Comparison Results
(a) Loss analysis
(b) Time cost comparison
(c) Different $d_{model}$
(d) Different window
Figure 5: The analysis of PatchAD and DCdetector.
4.3 Unveiling the Mechanisms
PatchAD leverages Inter and Intra Mixers to learn detailed and generalized features respectively, as demonstrated in Figure 8. Utilizing multiple patch sizes, as explored in Appendix (Figures 8(a) & 8(b)), enhances feature learning and generalization. The interplay between these mixers, visualized in Figure 3(c), effectively separates and amplifies distinct features, improving anomaly detection. Empirical validation, depicted in Figure 9, reveals that increasing patch size leads to more significant features acquired by the Intra Mixer, accelerating model learning. Finally, Figure 10 shows that anomalies are most easily detected in the Intra Mixer view due to its enhanced distinctiveness. In essence, maximizing the disparity between the Inter and Intra Mixers enhances PatchAD’s detection capabilities while mitigating overfitting. More details are provided in Appendix.
4.4 Model Analysis
4.4.1 Ablation Studies
Table 3: Ablation studies on MSL, SMAP, WADI datasets.
Dataset MSL SMAP WADI
Methods Acc P R F1 Acc P R F1 Acc P R F1
wo_ch_sharing 98.78 91.93 96.90 94.35 98.84 93.50 97.75 95.58 98.71 84.99 94.25 89.38
wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.50 97.10 95.27 98.92 85.40 97.93 91.24
wo_mixrep_mixer 96.41 89.54 74.66 81.43 95.25 91.52 68.55 78.70 95.76 67.55 51.99 58.76
wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74
wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53
PatchAD (Ours) 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43

### v6

@@ III-D Dual Project Constraint
Therefore, the outputs of the encoder can be represented as:
$\mathcal{N}=\alpha_{1}\cdot\mathcal{N}_{1}+\cdots+\alpha_{L}\cdot\mathcal{N}_{L},$ (13)
$\mathcal{P}=\alpha_{1}\cdot\mathcal{P}_{1}+\cdots+\alpha_{L}\cdot\mathcal{P}_{L}.$
@@
III-E Objective Function
The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$. To quantify the dissimilarity between these representations, we utilize a comparative loss function based on Kullback–Leibler divergence (KL divergence), termed as Inter-Intra Discrepancy. Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both. In Sec. IV-C1, the reasons for choosing the KL divergence as the basic loss are analyzed. The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
@@ III-F Anomaly Score
III-G Multi-Scale of PatchAD
For single-scale PatchAD, it features a comprehensive Patching, Patch Mixer Encoder, and Dual Project. This means that, assuming the patch size for single-scale PatchAD is $P_{i}$, we obtain the final output of the Patch Mixer Encoder, $\mathcal{\hat{X}}_{i}$, as well as the final outputs of the Dual Project, $\mathcal{N}_{i}$ and $\mathcal{P}_{i}$. These outputs are used to update the loss function $\mathcal{L}_{i}$ (as shown in Eq. (17)).
Here, $i$ denotes the $i$-th single scale. For a multi-scale PatchAD, there are typically $M$ branches. We optimize the loss functions $\{\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{M}\}$ in an integrated manner, allowing different branches to learn features at various scales. During the testing phase, each single-scale branch produces its own anomaly score $AS_{i}$, and the final anomaly score is the average of the scores from all branches, given by $AS=\frac{1}{M}\sum_{i=1}^{M}AS_{i}$.
IV Experiments
IV-A Setups and Evaluations
IV-A1 Datasets
We evaluated different algorithms using a total of 8 datasets, with details referred to in the appendix.
IV-A2 Metrics
Considering comprehensively, we adopt two groups of metrics to conduct extensive experiments. Detailed explanations for why these metrics are used can be found in Appendix -D3.
  1. 1.
    Metrics using point adjustment (PA): We prioritize these metrics because they are now the most commonly utilized group. To compare with advanced algorithms (including but not limited to [11, 17, 9, 46, 19, 15]) that used point-adjusted metrics for evaluation, we employ Acc, Pre, Rec, and PA-F1 as the main metrics.
  2. 2.
    Metrics without point adjustment: Meanwhile, as many current studies [28, 24] are gradually using classic metrics (AUC, F1-cls) and interval evaluation metrics (Aff-F1, VUS), in order to compare with these algorithms (including but not limited to [33, 47, 16, 48, 49, 50, 11]), we organize additional experiments for further analysis in Sec. IV-B2.
IV-A3 Baselines
We compared our model against 30+ benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) The classic methods; (6) Change point detection and time series segmentation methods; (7) Contrastive learning; (8) Large Language Model; (9) Diffusion model.
The details regarding the selection of the algorithms are further discussed in Appendix -D4.
IV-B Comparison Results
IV-B1 Comparisons via Metrics with PA
From Table I, PatchAD achieves the best average-level PA-F1 performance. PatchAD attains superior results across multiple datasets and multiple metrics. Despite PatchAD’s relatively simple network structure, its performance remains competitive. In the evaluation of the PA-F1 metric, DCdetector and AnomTrans emerge as the state-of-the-art(SOTA) models, with their robust capability to capture local anomalies granting them a competitive edge over other models in PA-F1. In contrast, PatchAD enhances its applicability to local anomalies through adaptive learning of local feature representations, achieving SOTA performance. Compared with AnomTrans’s limited temporal granularity, PatchAD’s multi-scale approach can better capture anomalies of different lengths, thus achieving superior PA-F1 performance.
Simultaneously, we observe that the additional models do not perform exceptionally well on the PA-F1 metric, primarily due to two reasons: firstly, they are mainly designed for the detection of interval anomalies, a fact that becomes evident in the subsequent analysis. Secondly, the PA-F1 metric employs point adjustment techniques, which can result in some undetected anomalies being classified as true positives, thereby introducing a potential inflation effect in the metric [51]. While PA-F1 is suitable for comparing the performance of different models in research contexts, it is less appropriate for assessing model performance in real-world scenarios. Consequently, this also accounts for the challenge faced by PatchAD in distinguishing itself from previous models. Therefore, in the following section, a new set of metrics will be employed to analyze the models from different perspectives.
@@ IV-B1 Comparisons via Metrics with PA
TABLE I: Comprehensive comparative results on four datasets. The P, R and PA-F1 are the precision, recall and F1 with PA. All results are in %, the best in Bold, and the second in underlined.
Dataset MSL SMAP PSM SWAT Average PA-F1
Metric P R PA-F1 P R PA-F1 P R PA-F1 P R PA-F1
LOF 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
OCSVM 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
U-Time 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
IForest 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
DAGMM 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
ITAD 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
VAR 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
MMPCACD 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
CL-MPPCA 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
TS-CP2 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
Deep-SVDD 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
BOCPD 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
LSTM-VAE 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
BeatGAN 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
LSTM 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
OmniAnomaly 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
InterFusion 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
THOC 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
AnomalyTrans 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
DCdetetor 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
TimesNet 89.54 75.36 81.84 90.14 56.40 69.39 98.51 96.20 97.34 90.75 95.40 93.02 85.40
PatchTST 88.34 70.96 78.70 90.64 55.46 68.82 98.84 93.47 96.08 91.10 80.94 85.72 82.33
GPT2-Adapter 82.00 82.91 82.45 90.60 60.95 72.88 98.62 95.68 97.13 92.20 96.34 94.23 86.67
@@
M2N2 [52] 89.38 92.48 90.90 87.37 72.15 79.04 70.51 89.99 79.07 99.42 70.96 82.82 82.96
AdaMemBLS [4] 88.01 93.20 90.53 88.01 93.20 90.53 89.59 95.45 92.43 93.19 82.78 87.68 90.29
PatchAD 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
@@
IV-B2 Comparisons via Metrics without PA
Figure 2 depicts the average performance of PatchAD compared to various baselines across multiple datasets (MSL, SMAP, SWAT, WADI, PSM, and SWAN). Here, Aff-F1 is the affiliation metric [22]. V-ROC and V-RR are volumes under the surfaces created based on ROC curve and PR curve [21], respectively. In terms of the F1-cls, Aff-F1, and V-ROC metrics, PatchAD demonstrates relative performance improvements of 6.84%, 4.27%, and 2.49% compared to the second-place model, respectively. This indicates a significant enhancement in PatchAD’s capacity for detecting interval anomalies compared to previous models. Notably, we observe that AnomTrans and DCdetector, which perform well on the PA-F1 metric, exhibit poor performance on interval anomaly metrics. This is primarily because these models are more focused on local or point anomalies, resulting in better performance on the PA-F1 metric. PatchAD, on the other hand, benefits from additional learning of generalized global features, enhancing its performance on interval anomalies, thus improving both PA-F1 and Aff-F1 metrics.
Furthermore, we note that although TimesNet was proposed earlier, it currently exhibits relatively strong performance when considering the two sets of metrics. This can be attributed to the diversity of contemporary anomalies, which vary in type, periodicity, and length. TimesNet’s capability to learn temporal information across different frequencies enables it to handle various anomalies more effectively. PatchAD similarly adopts this approach, with multi-scale temporal feature extraction ensuring the learning of information across different temporal frequencies.
NPSR also takes into account the utilization of two distinct feature types in its design and uniquely processes anomaly scores, thereby ensuring its excellent theoretical lower bound, which guarantees good performance in the face of real interval anomalies. Likewise, PatchAD benefits from the assurance provided by the lower bound of information entropy and the diverse features extracted through different Mixers, resulting in similarly outstanding performance. In contrast, COUTA’s training relies on anomaly injection, leading to varying adaptability across different real datasets; for instance, it performs well on the PSM and SWAN datasets but shows average performance on the WADI and SWAT datasets.
TABLE II: Comprehensive comparative results on six datasets. All results are in %, the best in Bold, and the second in underlined.
Datasets MSL SMAP SWAT
Methods AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR
LOF [2000] 55.75 19.36 66.86 60.85 17.19 62.39 23.64 58.74 57.56 18.59 84.59 58.22 5.25 66.67 41.88
IForest [2008] 59.09 10.90 70.44 64.74 18.29 61.15 23.78 50.52 59.31 16.45 83.80 38.52 5.27 72.29 52.07
PCA [2003] 53.25 10.33 70.72 59.42 18.08 58.70 22.96 50.72 43.90 11.69 81.85 26.05 5.25 61.87 44.75
Deep SVDD [2018] 61.80 24.30 68.98 59.54 19.36 61.14 21.28 59.64 41.95 11.56 86.81 22.57 69.77 50.96 56.48
USAD [2020] 57.12 22.50 67.69 58.86 19.77 62.79 16.57 50.41 33.78 11.28 88.66 21.70 69.28 84.38 62.74
TCN-ED [2021] 51.14 19.66 67.93 58.62 18.13 58.91 22.90 67.89 43.79 11.73 89.23 21.65 69.26 61.53 57.94
TranAD [2022] 50.03 22.81 68.97 57.25 17.55 59.74 22.74 50.51 45.15 11.98 88.90 25.50 71.42 86.25 64.19
NCAD [2022] 60.20 22.05 67.25 67.29 21.32 53.45 23.09 68.32 48.37 13.22 82.92 68.72 73.89 76.76 44.91
Deep IF [2023] 55.94 19.11 67.94 52.33 17.04 60.09 29.14 69.59 56.89 14.77 89.52 21.65 69.26 50.19 56.39
AnomTrans [2021] 52.61 18.39 66.81 53.03 13.21 52.18 16.06 71.44 52.61 16.55 80.80 23.44 66.77 21.49 8.29
TimesNet [2022] 57.18 21.24 67.64 64.63 20.67 53.73 24.37 66.19 48.83 12.91 73.24 21.66 71.55 31.10 10.29
GPT2-Ada. [2023] 52.03 13.72 68.81 53.57 17.38 55.48 24.12 69.48 57.31 15.78 52.30 22.30 68.41 50.01 10.29
DCdetector [2023] 50.31 11.62 67.85 52.00 14.68 58.50 26.56 71.37 60.38 16.87 52.78 23.24 68.56 50.84 14.68
D3R [2024] 63.00 23.98 69.79 69.02 20.93 54.56 22.71 67.92 48.67 13.23 79.95 45.89 68.96 84.38 57.27
COUTA [2024] 55.59 20.88 67.68 61.49 18.68 58.74 22.69 67.97 44.72 11.91 75.54 47.65 46.92 72.57 25.06
NPSR [2023] 61.16 23.72 68.42 66.44 19.72 38.74 22.68 67.95 41.91 11.42 90.18 76.88 75.86 84.93 65.42
M2N2 [2024] 59.15 21.76 67.88 65.40 21.55 54.05 22.92 67.95 48.74 14.36 67.79 38.50 73.44 61.14 26.16
AdaMemBLS [2025] 51.78 19.11 71.41 57.46 16.06 53.66 26.93 69.49 55.57 14.26 81.61 74.10 69.76 82.00 66.12
PatchAD 60.94 24.92 70.10 66.43 19.65 60.27 24.74 70.95 48.30 13.71 91.19 77.69 78.76 89.91 65.74
Datasets WADI PSM NIPS-TS-Swan
Methods AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR
LOF [2000] 65.44 10.58 34.36 66.67 41.88 81.28 23.91 62.32 80.42 64.55 77.89 21.37 19.62 89.42 88.56
IForest [2008] 74.84 30.88 67.90 72.29 52.07 75.12 38.52 56.77 75.19 55.50 85.12 46.18 27.81 91.27 90.98
PCA [2003] 49.47 34.30 50.01 61.87 44.75 72.06 43.20 78.74 73.97 54.55 72.88 49.18 27.35 90.42 90.23
Deep SVDD [2018] 59.47 7.34 52.77 50.96 56.48 78.09 44.46 70.06 53.97 66.92 75.32 55.99 69.66 80.89 91.74
USAD [2020] 53.61 10.86 70.90 84.38 62.74 64.53 47.90 70.88 64.02 69.18 66.76 52.63 61.22 91.40 94.50
TCN-ED [2021] 53.01 10.86 70.90 61.53 57.94 63.97 43.46 69.42 57.48 42.64 71.37 49.22 71.03 80.88 91.73
TranAD [2022] 52.60 11.78 69.54 86.25 64.19 65.22 43.20 73.30 72.42 56.02 70.00 51.93 70.49 88.01 92.59
NCAD [2022] 62.84 10.87 70.96 76.76 44.91 61.76 46.61 63.10 67.09 48.26 52.24 37.57 68.87 85.31 93.44
Deep IF [2023] 51.43 10.86 70.90 50.19 56.39 69.06 43.47 69.41 67.71 45.98 73.30 49.19 71.05 80.89 91.74
AnomTrans [2021] 52.98 11.24 66.77 21.49 8.29 52.18 39.81 68.13 49.12 31.56 54.62 44.86 69.11 83.56 81.64
TimesNet [2022] 65.08 17.65 49.05 31.10 10.29 58.74 43.60 72.15 66.37 48.05 53.62 43.16 69.53 91.58 94.30
GPT2-Ada. [2023] 51.21 10.68 71.04 50.01 10.29 51.23 35.24 69.17 52.45 32.79 58.77 45.23 65.09 84.31 84.23
DCdetector [2023] 50.12 11.33 73.23 50.84 14.68 50.38 22.72 68.08 51.73 32.44 50.35 48.83 70.99 85.84 83.11
D3R [2024] 51.39 12.86 72.17 84.38 57.27 60.64 44.24 69.99 66.92 49.38 62.54 49.45 72.10 89.27 88.29
COUTA [2024] 50.52 26.92 37.44 72.57 25.06 69.36 48.21 73.21 64.50 43.15 71.38 47.19 66.54 82.15 92.26
NPSR [2023] 80.19 50.67 75.74 84.93 65.42 70.67 51.02 70.56 70.55 52.12 62.61 49.30 71.07 90.90 90.35
M2N2 [2024] 55.68 13.55 70.90 43.99 7.98 54.91 47.71 69.84 40.06 30.61 70.27 57.47 71.05 89.77 90.18
AdaMemBLS [2025] 53.85 29.09 70.93 63.47 21.40 63.26 46.69 71.51 65.07 49.24 51.27 49.35 71.09 87.12 85.92
PatchAD 85.27 52.82 79.04 88.29 49.60 64.74 49.33 76.73 65.11 40.72 79.53 64.52 72.36 95.62 91.52
IV-C Model Analysis
IV-C1 Ablation Studies